# GPTVoiceTasker: LLM-Powered Virtual Assistant for Smartphone

### v1

@@ 3.3.4. Human Feedback Loop
The saved commands have proven their applicability in subsequent user interactions with the device, yet certain inconsistencies may arise due to app version updates and dynamic changes in UI element availability. To address error handling and corrections, GptVoiceTasker incorporates human feedback to enhance the saved data for each task. The automated tasks performed by the tool allow users to efficiently reverse actions and prompt the system to actively adjust the new execution path to accomplish tasks effectively. This valuable information is stored and fed back to the LLM, resulting in improved output for future executions. When users express satisfaction with the result and request the next task, we also increase the confidence score for this saved path, helping GptVoiceTasker to select the most reliable path to execute if multiple execution paths are found. In addition, we also provide an additional interface for users to amend saved commands manually. Through user feedback and iterative learning, GptVoiceTasker continuously refines its performance, ensuring increasingly accurate and effective responses over time.
3.4. Implementation
We implemented GptVoiceTasker as an Android application with the Accessibility Service in Android OS using Java (Developers, 2022b). Within the Java code, GptVoiceTasker subscribes to the typeWindowContentChanged accessibility events (Developers, 2022b) to receive notifications whenever UI changes occur on the screen. We created a dynamic pipeline to extract UI elements in a hierarchical structure from AccessibilityNodeInfo objects (Developers, 2022a), which serve as data representations of on-screen UI elements provided by Android Accessibility Service. In addition, We obtain the app name and the activity name using the Android PackageManager class.
To communicate with the LLM, we utilised the API service provided by OpenAI. We chose the GPT-4 model, as it is the latest model trained by OpenAI (OpenAI, 2023a). Once we receive a response from the LLM, the tool leverages the performAction() method (Developers, 2022a) to execute actions on the corresponding elements. All data related to personalized services, such as screen description and the transition graph, are stored in the phone memory for future usage.
@@ 4. Technical Evaluation
To evaluate the effectiveness and reliability of the proposed system, we conducted two experiments on our command interpreting module and usage-based execution. Specifically, we first assess the system’s ability to comprehend user commands and perform on-screen interactions, comparing its performance to other state-of-the-art approaches. We then investigate the system’s capability to execute multi-step tasks based on the saved user usages. .
@@ 4.1.1. Experiment Setup & Metric
Datasets: We collect a specialised test set to evaluate our system’s capabilities in understanding natural language commands and mapping them to appropriate actions and target UI elements. This dataset comprises 278 natural language user commands to interact with Android UIs.
@@
To construct a test set that more closely aligns with real-world user interactions, we adopted a data-driven approach. We engaged 31 participants (17 females, 14 males), with 4 individuals having never utilized voice assistants before, 4 using them 3-4 times a week, 6 using them daily, and 17 using them less than 3-4 times a week. All participants are work professionals and the rest from university community who use smartphones daily. We provided these participants with screenshots alongside a specific task to accomplish. We then recorded the verbal commands they issued to their mobile device to complete the given task. After the collection, we annotated the commands to specify the intended action and target UI elements within the Android system; here, the term action refers to executable functions, while target denotes specific UI components or elements on the current screen. As a result, we collected 278 natural user commands for the dataset.
@@
Baselines: We consider four baselines for converting natural language into semantic meaning representations, which comprise actions and targets. These baselines are vanilla Seq2Seq (Bahdanau et al., 2015), BERT-LSTM (Xu et al., 2020), Voicify Parser (Vu et al., 2023), and Wang et al.’s work (Wang et al., 2023a). In the original work by Voicify, all three baselines employ deep learning models trained on datasets synthesized using the Overnight method (Wang et al., 2015). This method generates training sets based on predefined lists of actions and targets that are designed for evaluation scenarios in Voicify. To ensure a fair comparison, we modified these lists to include the actions and targets present in our test dataset. We then re-synthesize the training set, which includes 1,384 instances, using the Overnight method, adhering to the implementation outlined in Voicify’s work. Wang et al.’s work (Wang et al., 2023a) was the first to incorporate LLMs for interacting with mobile interfaces. We incorporated the 2-shots LLM prompts they demonstrated in the paper for mapping instruction to UI actions. Additionally, we have enhanced their model by integrating the more advanced GPT-4, which also inline with the one we used in GptVoiceTasker.
4.1.2. Evaluation Result
Table 1 presents the results of our technical experiments. Overall, GptVoiceTasker significantly outperforms all baseline models across all metrics, achieving an 84.7% EM accuracy, a 91.7% Action F1 score, and a 84.7% Target F1 score. Among the baselines without the LLMs, the Voicify Parser performs the best, aligning with the results reported in its original paper (Vu et al., 2023). However, its performance suffers when faced with linguistic variations in our new test set. For instance, while the command “back” is correctly interpreted as “( PRESS , back )”, the phrase “return to last page”, which represents the same command, is incorrectly parsed as “( SWIPE , DOWN )”. Both BERT-LSTM and Seq2Seq models encounter similar issues, largely because they share architectural and training similarities with the Voicify Parser, yet perform even worse due to Voicify Parser being specifically optimized for task completion on Android systems. The method by Wang et al. (Wang et al., 2023a) demonstrates the highest capability among the baselines, courtesy of the LLM’s intervention, effectively rectifying the errors previously noted. However, the lack of the chain-of-thought and least-to-most prompt techniques occasionally leads to inaccuracies. This is evident in instances where the system misinterprets the intended direction in commands, such as confusing "DOWN" with "UP," or when it cannot adequately differentiate between actions like "PRESS," "ENTER," or "OPEN" when various verbs are employed in the commands.
Benefiting from the integration of LLM and the prompting techniques, our GptVoiceTasker excels at handling linguistic variants, consistently deriving the intended action and target regardless of variations in the input. The Action F1 score for GptVoiceTasker reaches 91.7, indicating its enhanced ability to predict actions across various linguistic patterns. Moreover, we observed that LLM effectively learns the true associations between actions and targets, thereby excelling at target prediction as well. For instance, PRESS is exclusively predicted with UI buttons, ENTER_TEXT is linked solely with text input fields, and OPEN corresponds to app names. In contrast, the baselines often learns incorrect associations and outputs wrong target predictions. Our experimental results demonstrate that GptVoiceTasker possesses superior capabilities for understanding and accurately processing various linguistic variations, highlighting its adaptability in real-world scenarios.
@@ 4.1.2. Evaluation Result
This table has 4 columns, first column represents the Model names as Sequence to sequence, BERT-LSTM, VoicifyParser and GPTVoiceTasker. The second to forth columns represent the exact match accuracy, Action F1 and Target F1. For Seq2seq, the exact match accuracy is 25.2%, action F1 is 47.6% and target F1 is 35.6%. For BERT-LSTM, the exact match accuracy is 41.4%, action F1 is 59.7% and target F1 is 57.3%. For VoicifyParser, the exact match accuracy is 47.5%, action F1 is 64.0% and target F1 is 58.8%. For GPTVoiceTasker, the exact match accuracy is 84.7%, action F1 is 91.7% and target F1 is 84.7%. Models EM Accuracy (%) Action F1 (%) Target F1 (%) Seq2Seq 25.2 47.6 35.6 BERT-LSTM 41.4 59.7 57.3 VoicifyParser 47.5 64.0 58.8 Wang et al. (Wang et al., 2023a) 79.9 85.4 83.4 GptVoiceTasker 84.7 91.7 84.7
@@ 4.2.1. Experimental Setup & Metric
In this experiment, we assessed GptVoiceTasker’s ability to automate tasks using the usage-based execution module. We initially identified the five common smartphone application categories, as shown in previous study (Arsan et al., 2021). Within each application category, we randomly selected five popular applications from the Google Play Store, with downloads ranging from 1 million to over 1 billion. For each selected app, we identified three features introduced by the developers in their Play Store descriptions. Each feature was then used to create both a direct match task, involving a straightforward match between user commands and corresponding app actions, and a parameterized task, requiring GptVoiceTasker to perform keyword substitutions to complete the task successfully, as shown in Section 3.3.3. For creating the direct match test cases, we paraphrased each saved command using state-of-the-art paraphrasing tool Quillbot [^3] https://quillbot.com/, as in (Shiri et al., 2022). In the case of parameterized tasks, we substituted one entity in the paraphrased command with another entity that has similar semantic. For example, consider the saved task “Get directions to the nearest supermarket”. In this case, the direct matching task would be “Find the nearest supermarket’s location”, while the parameterized task would involve substituting “restaurant” for “supermarket”, resulting in “Find the nearest restaurant’s location”. This process resulted in a total of five app categories, each category contains 15 direct match tasks and 15 parameterised tasks. These tasks involve 4 to 7 steps, with an average of 5.19 steps per task as illustrated in Table 2. All tasks can be automated with one voice command with the saved user app usage patterns. For a detailed list of the apps and features used in the experiment, please refer to our GitHub repository [^4] https://github.com/vuminhduc796/GPTVoiceTasker.
@@ 5. User Study
This table displays the list of tasks for user evaluation. Task 1 is ‘Check, send, and delete messages’ in messages app, with has 8 steps. Task 2 is ‘Search for a specific song and play it’ in Apple Music, with has 8 steps. Task 3 is ‘Create a new alarm and save it’ in Challenges Alarm Clock app, with has 10 steps. Task 4 is ‘Check the weather within a particular city’ in BOM Weather app, with has 6 steps. Task 5 is ‘Write and delete a note’ in Notes app, with has 8 steps. Task 6 is ‘Search for a pizza store, and complete the order’ in Uber Eats app, with has 10 steps. No. Task #Steps App Name #Downloads 1 Check the weather within a particular city. 6 BOM Weather 1M+ 2 Search for a specific song and play it. 6 Apple Music 100M+ 3 Create a note and write "Hello world" and delete it. 8 Notes 1M+ 4 Check for an unread message, reply with a message and delete the conversation. 8 Messages 1B+ 5 Search for a pizza store, and complete the order. 10 Uber Eats 100M+ 6 Create a new alarm and save it. 10 Challenges Alarm Clock 1M+
@@ 5.2. Participants
We recruited 18 participants, consisting of 10 males and 8 females, aged between 18 and 31 years old for our study. 8 participants are native English speaker while all other participants are proficient in English. All participants possess a commendable level of familiarity with technological devices and actively use smartphones in their daily routines. While participants exhibited exposure to virtual assistants like Siri or Google Assistant, none were acquainted with utilizing assistive tools for smartphone control via voice commands. Specifically, none of the participants had prior experience with any of the experimental tools employed in our study. This participant selection was deliberate, as our study sought to gauge the learnability aspect of the experimental tools. Each participant received a USD \$30 gift card for the participation.

### v2

@@ 3.2.4. Human Feedback Loop
One common issue with UI automation tools is dealing with changes in UI elements and in-app navigation due to new version updates by app developers. Additionally, pop-up ads may dynamically appear during the automation process. These challenges make predefined actions, such as clicking on an anticipated screen element becomes infeasible. To enhance the reliability of the saved paths, GptVoiceTasker incorporates a human-in-the-loop approach, modifying execution paths based on human feedback. After completing an autonomous task, GptVoiceTasker uses user feedback to validate whether the task was successfully executed. If users are unsatisfied with the automation, they can provide feedback in natural language, specifying where the issue occurs. This valuable information is saved and included in our prompts, resulting in improved accuracy for future executions. Additionally, we provide an interface for users to manually amend saved commands, allowing them to customize or shorten their commands to trigger certain automation according to their preferences. Through user feedback and iterative learning, GptVoiceTasker facilitates human-AI collaboration in the automation process, enhancing LLMs decision-making and personalizing user experiences.
$3.$ Implementation
We developed GptVoiceTasker as an Android application, leveraging the Accessibility Service provided by the Android OS and programming it in Java (Developers, 2022b). Specifically, GptVoiceTasker is designed to subscribe to the typeWindowContentChanged accessibility events (Developers, 2022b), enabling it to detect and respond to changes in the UI on the screen. We engineered a dynamic pipeline that systematically extracts UI elements and organizes them into a hierarchical structure. This is achieved by processing AccessibilityNodeInfo objects (Developers, 2022a), which are Android’s data representations of UI elements accessible through the Android Accessibility Service. Additionally, GptVoiceTasker gathers application-level information using the Android PackageManager class.
For the integration of a Large Language Model, we selected GPT-4, the most advanced model developed by OpenAI at the time of this research (OpenAI, 2023a). For personalized services, such as screen descriptions and transition graph data, we store this information locally on the device for efficient future access. We made GptVoiceTasker and the prompt templates [^2] https://github.com/vuminhduc796/GPTVoiceTasker/blob/main/prompts.txt publicly available at the GitHub repository [^3] https://github.com/vuminhduc796/GPTVoiceTasker for further research in this field.
@@ 4. Technical Evaluation
To evaluate the effectiveness and reliability of the proposed system, we conducted three experiments on our command interpreting module, unprecedented tasks exploration and usage-based execution. Specifically, we first assess the system’s ability to comprehend user commands and perform on-screen interactions, comparing its performance to other state-of-the-art approaches. In addition, we experiment the ability of GptVoiceTasker to explore the execution path for unseen multi-step tasks. Lastly, we investigate the system’s capability to execute multi-step tasks based on the saved user usages.
@@ 4.1.1. Experiment Setup & Metric
Datasets: We collect a specialised test set to evaluate our system’s capabilities in understanding natural language commands and mapping them to appropriate actions and target UI elements. This dataset comprises 278 natural language user commands to interact with on-screen Android UI elements.
@@
To construct a test set that more closely aligns with real-world user interactions, we adopted a data-driven approach. We engaged 31 participants (17 females, 14 males), with 4 individuals having never utilized voice assistants before, 4 using them 3-4 times a week, 6 using them daily, and 17 using them less than 3-4 times a week. All participants are work professionals and university students who use smartphones daily. We provided these participants with screenshots alongside a specific task to accomplish. We then recorded the verbal commands they issued to their mobile device to complete the given task. After the collection, we annotated the commands to specify the intended action and target UI elements within the Android system; here, the term action refers to executable functions, while target denotes specific UI components or elements on the current screen. As a result, we collected 278 natural user commands for the dataset.
@@
Baselines: We consider five baselines for converting natural language into semantic meaning representations, which comprise actions and targets. These baselines are vanilla Seq2Seq (Bahdanau et al., 2015), BERT-LSTM (Xu et al., 2020), Voicify Parser (Vu et al., 2023), the GptVoiceTasker w BasePrompt model, and Wang et al.’s work (Wang et al., 2023a). In the original work by Voicify, all three baselines employ deep learning models trained on datasets synthesized using the Overnight method (Wang et al., 2015). This method generates training sets based on predefined lists of actions and targets that are designed for evaluation scenarios in Voicify. To ensure a fair comparison, we modified these lists to include the actions and targets present in our test dataset. We then re-synthesize the training set, which includes 1,384 instances, using the Overnight method, adhering to the implementation outlined in Voicify’s work. The GptVoiceTasker with base prompt is used as an ablation study to the current prompt design of GptVoiceTasker. The base prompt contains only a minimal explanation of the task and the UI knowledge (XML file) of the current mobile screen. Wang et al.’s work (Wang et al., 2023a) was the first to incorporate LLMs for interacting with mobile interfaces. We incorporated the 2-shots LLM prompts they demonstrated in the paper for mapping instruction to UI actions. Additionally, we have enhanced their model by integrating the more advanced GPT-4, which also inline with the one we used in GptVoiceTasker.
@@
This table has 4 columns, first column represents the Model names as Sequence to sequence, BERT-LSTM, VoicifyParser and GPTVoiceTasker. The second to forth columns represent the exact match accuracy, Action F1 and Target F1. For Seq2seq, the exact match accuracy is 25.2%, action F1 is 47.6% and target F1 is 35.6%. For BERT-LSTM, the exact match accuracy is 41.4%, action F1 is 59.7% and target F1 is 57.3%. For VoicifyParser, the exact match accuracy is 47.5%, action F1 is 64.0% and target F1 is 58.8%. For GPTVoiceTasker, the exact match accuracy is 84.7%, action F1 is 91.7% and target F1 is 84.7%. Models EM Accuracy (%) Action F1 (%) Target F1 (%) Seq2Seq 25.2 47.6 35.6 BERT-LSTM 41.4 59.7 57.3 VoicifyParser 47.5 64.0 58.8 GptVoiceTasker w BasePrompt 58.4 71.6 62.4 Wang et al. (Wang et al., 2023a) 79.9 85.4 83.4 GptVoiceTasker 84.7 91.7 84.7
4.1.2. Evaluation Result
Table 1 presents the results of our technical experiments. Overall, GptVoiceTasker outperforms all baseline models across all metrics, achieving an 84.7% EM accuracy, a 91.7% Action F1 score, and a 84.7% Target F1 score. Among the baselines without the LLMs, the Voicify Parser performs the best, aligning with the results reported in its original paper (Vu et al., 2023). However, its performance suffers when faced with linguistic variations in our new test set. For instance, while the command “back” is correctly interpreted as “( PRESS , back )”, the phrase “return to last page”, which represents the same command, is incorrectly parsed as “( SWIPE , DOWN )”. Both BERT-LSTM and Seq2Seq models encounter similar issues, largely because they share architectural and training similarities with the Voicify Parser, yet perform even worse due to Voicify Parser being specifically optimized for task completion on Android systems. Despite the lack of prompt design, GptVoiceTasker with a base prompt achieved better performance than the other three DL-based baselines. However, when compared with GptVoiceTasker, the results indicate that incorporating a multi-level knowledge-based prompt design, along with few-shot learning and the Chain of Thought technique within GptVoiceTasker, can enhance the accuracy of converting natural language to on-screen actions and the corresponding elements. The method by Wang et al. (Wang et al., 2023a) demonstrates the highest capability among the baselines, courtesy of the LLM’s intervention, effectively rectifying the errors previously noted. However, the lack of the chain-of-thought and least-to-most prompt techniques occasionally leads to inaccuracies. This is evident in instances where the system misinterprets the intended direction in commands, such as confusing DOWN with UP, or when it cannot adequately differentiate between actions like PRESS ENTER or OPEN when various verbs are employed in the commands.
Benefiting from the integration of LLMs and the prompting techniques, GptVoiceTasker excels at handling linguistic variants, consistently deriving the intended action and target regardless of variations in the input. The Action F1 score for GptVoiceTasker reaches 91.7, indicating its enhanced ability to predict actions across various linguistic patterns. Moreover, we observed that LLMs effectively learns the true associations between actions and targets, thereby excelling at target prediction as well. For instance, PRESS is exclusively predicted with UI buttons, ENTER_TEXT is linked solely with text input fields, and OPEN corresponds to app names. In contrast, the baselines often learns incorrect associations and outputs wrong target predictions. Our experimental results show that GptVoiceTasker is effective in understanding and accurately processing different linguistic variations, demonstrating its adaptability in real-world scenarios.
4.2. Multi-step Execution Evaluation
4.2.1. Experimental Setup & Metrics
This experiment assesses the Unprecedented Task Exploration module’s capability to execute unseen tasks. We evaluated the module’s performance using the most recent human-collected demonstrations and natural language instructions from the Android-in-the-wild dataset (Rawles et al., 2024). This dataset provides step-by-step on-screen interactions to complete tasks based on natural language instructions, mirroring real-world commands. From this dataset, we randomly sample the data from multi-step Google Apps subset and manually validate each pair to get 140 test cases, with each test case has from 1 to 15 steps (M=6.705, SD=2.764). We treat the end screen after the final action in the action sequence from the dataset as the ground truth, indicating successful command execution. We did not evaluate the accuracy of each steps to the dataset as one task could be performed by multiple approach. For each test case, we have GptVoiceTasker iteratively explore the path to complete the instruction, comparing the destination screen achieved by GptVoiceTasker with the dataset’s ground truth. A test case was deemed successful if GptVoiceTasker reached the same screen as it is in the ground truth with no more than three additional steps than the dataset demonstration. Cases where the step count was exceeded or the next step could not be identified were marked unsuccessful. GptVoiceTasker executed commands solely using its task exploration module, without relying on a database for guidance. Similar to Sec 4.1, we used Wang et al. (Wang et al., 2023a) approach as the baseline, which used a 2-shot prompting technique on the GPT-4 model, making iterative requests after each action response.
4.2.2. Results
GptVoiceTasker achieved an 85.7% success rate (120 out of 140), outperforming the baseline approach by Wang et al., which achieved a 56.4% success rate (79 out of 140). GptVoiceTasker succeeded in performing logical reasoning to execute the tasks. It utilizes the current app name and package name to determine if required apps need to be opened, and uses the list of run-time device-available app names to query the most suitable app for the task. Notably, GptVoiceTasker effectively navigated tasks without encountering cyclic navigation issues or repeating actions, such as repeatedly tapping the Settings title within the Settings app, which hindered task completion in the baseline. This improvement is attributed to the integration of historical messages and runtime execution error handling module. Our study also found that GptVoiceTasker does not rigidly adhere to the demonstrated paths presented in the Android-in-the-wild dataset for task completion. For example, in managing tasks within the Android Settings App, it occasionally opted to search directly for options rather than scrolling through menus to locate them. Other added constraints in GptVoiceTasker helped to improve the usability of multi-step interactions, such as validating if the current screen is scrollable or a text field is focused before inserting text, which were observed as reasons that caused failures in the baseline.
Further investigation into unsuccessful cases highlighted two main areas for improvement. First, GptVoiceTasker struggled with time-related tasks, such as “Check the schedule for Friday next week”, due to the LLM’s limited knowledge of the current date and time. As a result, while GptVoiceTasker correctly opened the app, it was unable to select the Friday of next week to view the schedule. This could be mitigated by integrating temporal information into our prompts. Second, GptVoiceTasker exceeded the number of steps when performing “Open the settings page in Google Maps”, where the immediate step required pressing the profile picture, which is not directly relevant to the task command. To address this issue, we could enrich the prompts with additional contextual knowledge, such as the steps involved in accessing settings in similar applications where pressing the profile picture is necessary. Additionally, incorporating a strategy of random exploration steps before repeating a command could help the system discover more direct paths to complete tasks.
@@ 4.3.1. Experimental Setup & Metric
In this experiment, we assessed GptVoiceTasker’s ability to automate tasks using the usage-based execution module. We initially identified the five common smartphone application categories, as shown in previous study (Arsan et al., 2021). Within each application category, we randomly selected five popular applications from the Google Play Store, with downloads ranging from 1 million to over 1 billion. For each selected app, we identified three features introduced by the developers in their Play Store descriptions. Each feature was then used to create both a direct match task, involving a straightforward match between user commands and corresponding app actions, and a parameterized task, requiring GptVoiceTasker to perform keyword substitutions to complete the task successfully, as shown in Section 3.2.3. For creating the direct match test cases, we paraphrased each saved command using state-of-the-art paraphrasing tool Quillbot [^4] https://quillbot.com/, as in (Shiri et al., 2022). In the case of parameterized tasks, we substituted one entity in the paraphrased command with another entity that has similar semantic. For example, consider the saved task “Get directions to the nearest supermarket”. In this case, the direct matching task would be “Find the nearest supermarket’s location”, while the parameterized task would involve substituting “restaurant” for “supermarket”, resulting in “Find the nearest restaurant’s location”. This process resulted in a total of five app categories, each category contains 15 direct match tasks and 15 parameterised tasks. These tasks involve 4 to 7 steps, with an average of 5.19 steps per task as illustrated in Table 2. All tasks can be automated with one voice command with the saved user app usage patterns. For a detailed list of the apps and features used in the experiment, please refer to our GitHub repository [^5] https://github.com/vuminhduc796/GPTVoiceTasker/blob/main/Result.xlsx.
@@ 5. User Study
This table displays the list of tasks for user evaluation. Task 1 is ‘Check, send, and delete messages’ in messages app, with has 8 steps. Task 2 is ‘Search for a specific song and play it’ in Apple Music, with has 8 steps. Task 3 is ‘Create a new alarm and save it’ in Challenges Alarm Clock app, with has 10 steps. Task 4 is ‘Check the weather within a particular city’ in BOM Weather app, with has 6 steps. Task 5 is ‘Write and delete a note’ in Notes app, with has 8 steps. Task 6 is ‘Search for a pizza store, and complete the order’ in Uber Eats app, with has 10 steps. No. Task #Steps App Name #Downloads 1 Check the weather within a particular city. 6 BOM Weather 1M+ 2 Search for a specific song and play it. 6 Apple Music 100M+ 3 Create a note and write ”Hello world” and delete it. 8 Notes 1M+ 4 Check for an unread message, reply with a message and delete the conversation. 8 Messages 1B+ 5 Search for a pizza store, and complete the order. 10 Uber Eats 100M+ 6 Create a new alarm and save it. 10 Challenges Alarm Clock 1M+
@@ 5.2. Participants
We recruited 18 participants, consisting of 10 males and 8 females, aged between 18 and 31 years old for our study. The mean age of participants was 25.83 years (SD = 4.26). The group included 8 bachelor students (from IT and Business fields), 4 master students (IT), and 6 PhD candidates. 8 participants are native English speaker while all other participants are proficient in English. All participants possess a commendable level of familiarity with technological devices and actively use smartphones in their daily routines.
We advertised our experiment on LinkedIn to recruit participants from our university. During recruitment, participants provided their gender, age, study level, English proficiency, tech proficiency, and experience with voice assistants. While participants exhibited exposure to virtual assistants like Siri or Google Assistant, none were acquainted with utilizing assistive tools for smartphone control via voice commands. Specifically, none of the participants had prior experience with any of the experimental tools employed in our study. This participant selection was deliberate, as our study sought to gauge the learnability aspect of the experimental tools. Each participant received a USD \$30 gift card for the participation.